OpenStudy (anonymous):

Change these polar equations to rectangular form θ = 135°

zepdrix (zepdrix):

\[\Large\bf\sf \arctan\left(\frac{y}{x}\right)\quad=\quad \theta\qquad\implies\qquad \tan 135^o\quad=\quad \frac{y}{x}\] \[\Large\bf\sf y=(\tan 135^o) x\]

zepdrix (zepdrix):

Hmm that's kind of confusing to do it that way ^ lemme draw a graph instead.

OpenStudy (anonymous):

So your good at these? i have one more

zepdrix (zepdrix):